Problem 8:




A vertically downward force of magnitude 10 N is required to immerse a homogeneous object of weight 40 N completely in water. What is the density of the object?

1
Expert's answer
2022-04-11T12:03:28-0400

"\rho_w g\frac m\rho =P+F,"

"\frac {\rho _w}{\rho}P=P+F,"

"\frac{\rho_w}{\rho}=\frac{P+F}P,"

"\rho=\rho_w\frac P{P+F},"

"\rho=800~\frac {kg}{m^3}."